40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Feb 2 18:35:47 2016, gautam, Update, Green Locking, Lightwave NPRO moved from PSL table to SP table 
    Reply  Tue Feb 2 18:54:02 2016, Koji, Update, Green Locking, Lightwave NPRO moved from PSL table to SP table 
       Reply  Wed Feb 3 23:23:47 2016, gautam, Update, Green Locking, Lightwave NPRO moved from PSL table to SP table KnifeEdgeScans_x.pdfBeamscan_x.pdf
          Reply  Thu Feb 4 09:16:46 2016, Koji, Update, Green Locking, Lightwave NPRO moved from PSL table to SP table 
             Reply  Fri Feb 5 00:23:01 2016, gautam, Update, Green Locking, Lightwave NPRO moved from PSL table to SP table Beamscan_setup.pdfBeamscan_x.pdfBeamscan_y.pdfZscan.pdf
                Reply  Fri Feb 5 16:50:24 2016, gautam, Update, Green Locking, First pass at mode-matching Modematch_AUXx.pdfNewSetUp.pngModematch_AUXx_2.pdfXendModeMatch.m.zip
                   Reply  Mon Feb 8 15:36:37 2016, gautam, Update, Green Locking, Alternative mode-matching scheme Modematch_alternative.pdf
                      Reply  Tue Feb 9 04:37:10 2016, ericq, Update, Green Locking, Laser swap initiated 
                         Reply  Tue Feb 9 11:49:47 2016, gautam, Update, Green Locking, Laser swap initiated Innolight_beamProfile.pdf
                            Reply  Tue Feb 9 19:15:36 2016, gautam, Update, Green Locking, Laser swap - updates Modematch_X.pdfXendModeMatch.m.zip
                               Reply  Wed Feb 10 17:57:15 2016, gautam, Update, Green Locking, Laser swap - updates 
                                  Reply  Fri Feb 12 17:05:40 2016, gautam, Update, Green Locking, Laser swap - green light recovered but no flashes in the arm 
                                     Reply  Fri Feb 12 19:07:52 2016, Koji, Update, Green Locking, Laser swap - green light recovered but no flashes in the arm 
                                        Reply  Fri Feb 19 00:42:19 2016, gautam, Update, Green Locking, Laser swap - beam ellipticity from laser? IMG_6287.JPGIMG_6288.JPGIMG_6286.JPG
                                           Reply  Mon Feb 22 13:56:52 2016, gautam, Update, Green Locking, Laser swap -reflected beam from ETM aligned 
                                              Reply  Tue Feb 23 23:01:16 2016, gautam, Update, Green Locking, Laser swap - Green PDH locking AM_scan.pdfAM_scan_zoomed.pdf
                                                 Reply  Wed Feb 24 19:29:13 2016, gautam, Update, Green Locking, Laser swap - Green PDH locked 
                                                    Reply  Fri Feb 26 01:52:44 2016, gautam, Update, Green Locking, Laser swap - Green PDH OLTF X_PDH_OLTF_20160225.pdf
                                                       Reply  Mon Feb 29 17:17:26 2016, gautam, Update, Green Locking, Laser swap - still no green beatnote 
                                                          Reply  Wed Mar 2 17:42:19 2016, gautam, Update, Green Locking, Laser swap - some progress 
                                                             Reply  Thu Mar 3 01:25:50 2016, gautam, Update, Green Locking, Laser swap - 2 IR + 1 green beatnotes found IR_beat_20160303_2.pdf
                                                                Reply  Fri Mar 4 01:11:41 2016, gautam, Update, Green Locking, Laser swap - both green beatnotes found IR_beat_20160303_green.pdf
                                                                   Reply  Sat Mar 5 23:31:01 2016, gautam, Update, Green Locking, Laser swap - some updates AUXxTelescope.png.pngPM_setup.pdfPLLolg.pdfAMPM20160303.pdfIRbeat_20160304.pdf
                                                                      Reply  Mon Mar 7 23:51:36 2016, gautam, Update, Green Locking, Laser swap - some improvement PomonaTF.pdfXPDH.pdfYPDH.pdfgreenbeat_20160307.pdf
                                                                         Reply  Tue Mar 8 18:22:20 2016, rana, Update, Green Locking, Laser swap - some improvement 
Message ID: 12023     Entry time: Sat Mar 5 23:31:01 2016     In reply to: 12019     Reply to this: 12026
Author: gautam 
Type: Update 
Category: Green Locking 
Subject: Laser swap - some updates 

I've been a little behind on my elogs so here is an update of the end laser situation.

IR beat for X-end recovered

  • The issue was optimizing the alignment into the fiber at the end table.
  • Using Fluke fiber illuminator helped in aligning IR pickoff into mount. Useful note: there is an unused fiber running between the X-end and the PSL table, by connecting these at the PSL table, I was able to monitor the coupled power while remaining at the X-end.
  • Another major issue was that one of the steering mirrors (marked "Y1" in Attachment #1) was mounted with AR coated side facing the beam. This was fixed by simply rotating the post, the mirror was not removed from its mount. I can only assume that this mirror is in this kind of mount because of space constraints.
  • The fiber has a collimating telescope attached to the end of it. In principle, this gives us more angular acceptance while coupling the beam into the fiber, but as I found out, the acceptance is still tiny (I don't have a number to quantify it). Furthermore, the Fluke visual fault locator revealed that the lens in the collimating telescope is not set up great - when re-doing the X end table, we should fix this situation so as to have a fairly large collimated beam coming out of the fiber when illuminated from the other end, this would make the mode matching much easier.
  • Bottom line: we have ~1.2 mW of IR light incident on the coupler at the end table, and ~400uW of IR power at the PSL table => coupling efficiency is ~30%, not stellar, but sufficient for now I guess. After the various splitters etc, there is about 160uW of EX IR light and ~300uW of PSL IR light incident on the beat PD, and the beat amplitude is about -9dBm.

AM/PM characterization of newly installed Lightwave

  • Having recovered the IR beat, I set out to do the PM characterization for the end laser.
  • Attachment #2 shows the electrical setup. The IR beat was piped to the X-end via an existing long cable that runs between the vertex and the endtable. Not shown in the diagram, but I used a 20dB coupler to keep track of the beat frequency on the HP spectrum analyzer while doing this measurement.
  • I restricted myself to the range between 100kHz and 500kHz to do the scan, because it takes quite a while to do the scan with fine resolution (IF bandwidth = 10Hz).
  • To calibrate the magnitude response to rad/V, I divided the output of the network analyzer (converting dB to absolute magnitude first) by the amplitude of the signal seen on the monitoring oscilloscope while the PLL is unlockedThis number was 96mV/rad.
  • To confirm that the error signal spectrum is indeed a good approximation of the "plant" transfer function (i.e that 100kHz >> UGF of loop transfer function of the PLL), I measured the loop TF of the PLL - Attachment #3 suggests a UGF of ~ 16kHz, which means the assumption is reasonable.
  • Excitation amplitude was -25dBm (which gave reasonable SNR), and 3 averages were taken.
  • The AM measurement was done using the same procedure as detailed here - the DC block was used. The DC level of the PD output was 2.72 V. The excitation amplitude was 0dBm.
  • Attachment #4 shows the AM response, PM response and PM/AM ratio
  • The peak in the PM/AM ratio at 256620 Hz is compelling because it is not too sharp (and so we can be reasonably confident we are at a good operating point) and the PM response of 23.83 rad/V is also acceptable. 
  • As a consistency check, the PM response of ~30rad/V at 100kHz => PZT actuator gain is ~3MHz/V, which is in the region we expect it to be...

Next steps in recovering ALS and trying to lock again

  • Having set the PDH modulation frequency to 256.62kHz, I took the spectrum of ALS noise using the IR beat (i.e. by piping the IR beat signal through the electronics the green beats usually go through - 6dB and 10dB attenuators were placed immediately after the beat PDs for the X and Y arms respectively, to make the signal levels compatible with the electronics), Attachment #5 unfortunately suggests that the noise performance is still poor, and I suspect the situation will be similar using the green beat (though I have not measured this yet).
  • The modulation depth could be sub-optimal for the X-end PDH, I have to measure this and check that it is at an acceptable level. This will also tell me if I need to change the sum+HPF pomona box used to send the PDH control signal + piezo dither signal to the laser PZT. In order to do this, I need to know what the input impedance to the FAST control BNC is - the manual isn't very helpful, it just says the piezo has a capacitance less than 10,000pF. I suppose I will have to actually measure this.
  • PDH loop OLTFs have to be re-measured for both ends to check that the servo gain's are appropriately placed.
  • We know that the mode-matching into the arm for the X end is poor (I have yet to quantify this) - I suspect that the beam ellipticity is the main culprit. However, the DC transmitted power levels at the PSL table are comparable to (even slightly better than) the Y arm numbers, and so this cannot be the sole reason why the X-arm ALS noise is so much worse... I will continue my investigations next week...
Attachment 1: AUXxTelescope.png.png  662 kB  | Hide | Hide all
AUXxTelescope.png.png
Attachment 2: PM_setup.pdf  119 kB  | Hide | Hide all
PM_setup.pdf
Attachment 3: PLLolg.pdf  48 kB  | Hide | Hide all
PLLolg.pdf
Attachment 4: AMPM20160303.pdf  62 kB  | Hide | Hide all
AMPM20160303.pdf
Attachment 5: IRbeat_20160304.pdf  62 kB  | Hide | Hide all
IRbeat_20160304.pdf
ELOG V3.1.3-